OpenStudy (anonymous):

BIG ASSIGNMENT DUE TOM!!! How do you find the last point of an equilateral triangle on a coordinate grid?

OpenStudy (anonymous):

Please... help... me.... I'm desperate...

OpenStudy (anonymous):

Will attempt to draw a picture of your choosing if you successfully help me!

OpenStudy (anonymous):

well equilateral triangles all have equal sides lengths. find the distance between the first two points and the last point will be the same distance away from your first two points

OpenStudy (anonymous):

I have been furiously working scribbling onto my notepad and have arrived at the conclusion that \[\sqrt{(1-0)^{2}}+(X-0)^{2}=\sqrt{(1-0)^{2}}+(X-0)^{2}\] although that is nice and all that ends up with 1=1.